The enthalpies of combustion of carbon and carbon monoxide are -393.5
and -283 kJ mol -1 respectively. The enthalpy of formation of carbon
monoxide per mole is?
 

Grade:12

3 Answers

Sakshi
8 years ago
hey student..
for this question you need to keep in mind the definations of both the enthalipies written over there in the question. then try writing the reactions one by one. also try to use the hess law of thermodynamics for the above mentioned question and you ll be getting the answer..
arun
8 years ago
in this given enthalpy of combustion for C then reaction for combustion of carbon  C_{(s)} + O_{2}_{(g)} \rightarrow CO_{2}_{(g)}          \Delta H= - 393.5 KJ/mol
and reaction for combustion of carbon-mono-oxide is CO_{(g)} + \frac{1}{2}O_{2(g)} \rightarrow CO_{2(g)}  \Delta H= -283KJ/mol
reversing this equation CO_{2(g)} \rightarrow CO_{(g)} + \frac{1}{2}O_{2(g)}   \Delta H = 283 KJ/mol
adding this equation with the equation of enthalpy of combustion of carbon
C_{(s)} + \frac{1}{2}O_{2(g)} \rightarrow CO_{(g)}      \Delta H = -393.5 + 283 = 110.5 KJ/mol
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Write both the equation with their enthalpy of combustion.
Write the desired equation, and use the hess law to obtain your answer,
We get, ΔH = – 393.5 + 283 = 110.5 KJ/mol
 
Thanks and regards,
Kushagra